OpenStudy (anonymous):

Rationalize the denominator of square root of negative 49 over open parentheses 7 minus 2 I close parentheses minus open parentheses 4 plus 9 i

OpenStudy (richyw):

\[\frac{-49}{\left(7-2i\right)\left(4+9i\right)}\]

OpenStudy (richyw):

is that what you want? please use the equation editor or type the equations using LaTeX code.